A monopoly firm is faced with the following demand function p=13_0.5Q the marginal cost function for the firm is given by 3+4Q and the total fixed cost is four


Determine


(a) The profit maximizing output


(b) The level of supernormal output if any


(C) the output level at the break even point

1
Expert's answer
2022-01-24T10:48:08-0500

"P= 13-0.5Q"

TR=P.Q

"TR= 13Q-0.5Q^2"

MR=13-Q

MC=3Q+4Q

TC=3Q+2Q"^2"


1) For profit Maximization,

MR=MC

"13-Q=3+4Q"

5Q=10

Q= 2

P=13-0.5(2)= 12

2) Supernormal output

Supernormal print occurs when average revenue is greater than Average total cost

AR="\frac{13Q-0.5Q^2}Q= 13-0.5Q"

=13-0.5(2)= 12

Ac= "\frac{3Q+2Q^2}{Q}= 3+2Q"

=3+2(2)=7

AR is greater than AC by 12-7= 5

3) Break even output

TR=TC

"3Q+2Q^2=13Q-0.5Q^2"

"2.5Q^2-10Q=0"

"Q(2.5Q-10)=0"

Q=0 0r

2.5Q=10

Q= "\frac{10}{2.5}= 4"